### v2

# MV-VTON: Multi-View Virtual Try-On with Diffusion Models

Abstract.
The goal of image-based virtual try-on is to generate an image of the target person naturally wearing the given clothing. However, most existing methods solely focus on the frontal try-on using the frontal clothing. When the views of the clothing and person are significantly inconsistent, particularly when the person’s view is non-frontal, the results are unsatisfactory. To address this challenge, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to reconstruct the dressing results of a person from multiple views using the given clothes. On the one hand, given that single-view clothes provide insufficient information for MV-VTON, we instead employ two images, i.e., the frontal and back views of the clothing, to encompass the complete view as much as possible. On the other hand, the diffusion models that have demonstrated superior abilities are adopted to perform our MV-VTON. In particular, we propose a view-adaptive selection method where hard-selection and soft-selection are applied to the global and local clothing feature extraction, respectively. This ensures that the clothing features are roughly fit to the person’s view. Subsequently, we suggest a joint attention block to align and fuse clothing features with person features. Additionally, we collect a MV-VTON dataset, i.e., Multi-View Garment (MVG), in which each person has multiple photos with diverse views and poses. Experiments show that the proposed method not only achieves state-of-the-art results on MV-VTON task using our MVG dataset, but also has superiority on frontal-view virtual try-on task using VITON-HD and DressCode datasets. Codes and datasets will be publicly released at https://github.com/hywang2002/MV-VTON.
Virtual try-on, Multi-view, Diffusion models
ccs: Computing methodologies Computer vision
1. Introduction
Virtual Try-On (VTON) is a classic yet intriguing technology. It can be applied in the field of fashion, and clothes online shopping to improve user experience. VTON aims to render the visual effect of a person wearing a specified garment. The emphasis of this technology lies in reconstructing a realistic image that faithfully preserves personal attributes (e.g., appearance and pose) and accurately represents clothing shape and details.
Most early VTON methods (Choi et al., 2021; Lee et al., 2022; Xie et al., 2023; Bai et al., 2022; Han et al., 2019; He et al., 2022; Li et al., 2021; Yang et al., 2020) are based on generative adversarial networks (Goodfellow et al., 2020) (GANs). They generally align the clothing to the person’s pose through a warping network, and then employ a generator to fuse the warped clothing with the person. However, it poses a challenge to ensure that the warped clothing fits the target person’s pose, and inaccurate clothes features will easily lead to distortion results. Recently, the diffusion models (Rombach et al., 2022; Zhang et al., 2023; Huang et al., 2023c; Nichol and Dhariwal, 2021) have made remarkable strides in the field of image generation (Li et al., 2023; Ruiz et al., 2023; Huang et al., 2023a; Xu et al., 2023; Huang et al., 2023b). Leveraging its potent generative capabilities, some researchers (Morelli et al., 2023; Gou et al., 2023; Kim et al., 2023; Zhu et al., 2023; Xu et al., 2024) have integrated it into virtual try-on endeavors, building upon previous work and achieving commendable results. For example, LaDI-VTON (Morelli et al., 2023) encodes the garment through textual inversion as the conditional input for Stable Diffusion (Rombach et al., 2022) backbone, and introduces a residual connection between the VAE (Kingma and Welling, 2013) encoder and decoder to preserve person’s face and hands. DCI-VTON (Gou et al., 2023) attaches the pre-warped garment to the inpainting area, providing a priori information for the diffusion model. StableVITON (Kim et al., 2023) introduces an additional encoder and attention blocks to deform the clothing.
Although VTON has made great progress, most existing methods focus on taking the frontal garment to perform the frontal try-on. In practical applications, such as online shopping for clothes, customers may expect to obtain the dressing effect on multiple views (e.g., side or back). In this case, the pose of the garment may be seriously inconsistent with the person’s posture, and the single-view clothing may not be enough to provide complete try-on information. Thus, these methods will easily generate results with poorly deformed clothing, and lead to the loss of high-frequency details such as texts, patterns, and other textures on clothing.
To address the issue, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to faithfully reconstruct the appearance and attire of a person from various views. MV-VTON should not only address the challenge of natural warping on the clothing’s front but also tackle the issue of inconsistency between the poses of the given clothing and target person. For example, for hooded sweaters, which may exhibit significant differences between frontal and back styles, MV-VTON should display them from various views to provide a more comprehensive fitting result. Thus, only providing single clothing can’t meet the needs of dressing up. Instead, we utilize both the frontal and back views of the clothing, which covers approximately the complete view with as few images as possible.
Given the frontal and back clothing, we utilize the popular diffusion method to achieve MV-VTON. It is natural but doesn’t work well to simply concatenate two pieces of clothing together as conditions of diffusion models, as it is difficult for the model to learn how to assign two-view clothes to a person, especially when the person is sideways. Instead, we propose a view-adaptive selection mechanism, which picks appropriate features of two-view clothes based on the posture information of the person and clothes. Therein, the hard-selection module chooses one of the two clothes for global feature extraction, and the soft-selection module modulates the local features of two clothes. We utilize CLIP (Radford et al., 2021) and an additional multi-scale encoder to extract the global and local clothing features, respectively. Moreover, to enhance the preservation of high-frequency details in clothing, we present the joint attention block. It independently aligns global and local features with the person features, and selectively fuses them to refine the local clothing details while preserving global semantic information.
Furthermore, we collect a multi-view virtual try-on dataset, named Multi-View Garment (MVG). It contains thousands of samples, and each sample contains 5 images under different views and poses. We conduct extensive experiments not only on MV-VTON task using the MVG dataset, but also on the frontal-view VTON task using VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. The results demonstrate that our method outperforms existing methods on both tasks, quantitatively and qualitatively.
In summary, our contributions are outlined below:
  • We introduce a novel Multi-View Virtual Try-ON (MV-VTON) task, which aims at generating realistic dressing-up results of the multi-view person by using the given frontal and back clothing.
  • We propose a view-adaptive selection method, where hard-selection and soft-selection are applied to global and local clothing feature extraction, respectively. It ensures that the clothing features are roughly fit to the person’s view.
  • We propose the joint attention block to align the global and local features of selected clothing with the person ones, and fuse them.
  • We collect a multi-view virtual try-on dataset. Extensive experiments demonstrate that our method outperforms previous approaches quantitatively and qualitatively in both frontal-view and multi-view virtual try-on tasks.
2. Related Work
2.1. GAN-Based Virtual Try-On
Most of the previous methods are aimed at the frontal-view VTON task. To reconstruct realistic results, existing methods based on generative adversarial networks (GAN) (Goodfellow et al., 2020) are typically divided into two steps. Firstly, the frontal-view clothing is deformed to initially align with the target person’s pose. Afterward, the warped clothing and target person are fused through a GAN-based generator. In the warping step, some methods (Morelli et al., 2023; Yang et al., 2020; Ge et al., 2021a; Wang et al., 2018) use TPS transformation to deform the frontal-view flattened clothing, and others (Lee et al., 2022; Ge et al., 2021b; Xie et al., 2023) predict the global and local optical flow required for clothing deformation. However, when the clothing possesses intricate high-frequency details and the person’s pose is complex, the effectiveness of clothing deformation is often significantly diminished. Moreover, GAN-based generators generally encounter challenges in convergence and are highly susceptible to mode collapse (Miyato et al., 2018), leading to noticeable artifacts at the junction between warped clothing and the target person in the final results. In addition, previous multi-pose virtual try-on methods (Dong et al., 2019; Wang et al., 2020; Yu et al., 2023) can change the person’s pose, but are also limited by GAN-based generator and insufficient clothing information. In this paper, we leverage the potent generative capability of a large-scale pre-trained diffusion model to implicitly deform the clothing from frontal and back views. Simultaneously, we employ the view-adaptive selection to choose the clothing features from the two views.
2.2. Diffusion-Based Virtual Try-On
Thanks to the rapid advancement of diffusion models, recent works have sought to utilize the generative prior of large-scale pre-trained diffusion models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022; Yang et al., 2023) to tackle frontal-view virtual try-on tasks. TryOnDiffusion (Zhu et al., 2023) introduces two U-Nets to encode target person and frontal-view clothing images respectively, and interacts with the features of the two branches through the cross-attention mechanism. LaDI-VTON (Morelli et al., 2023) encodes the frontal-view clothing image through textual inversion (Gal et al., 2022; Wei et al., 2023) and serves as the conditional input of backbone. DCI-VTON (Gou et al., 2023) first conducts an initial deformation of frontal-view clothing by incorporating a pre-trained wrapping network (Ge et al., 2021b). Subsequently, it attaches the deformed clothing to the target person image and feeds it into the diffusion model. While their frontal-view virtual try-on results seem more natural compared to GAN-based methods, they face difficulties in preserving high-frequency details due to the loss of details from the CLIP image encoder (Radford et al., 2021). To address this problem, StableVITON (Kim et al., 2023) attempts to introduce an additional encoder (Zhang et al., 2023) to encode the local features of frontal-view clothing, and align the obtained clothing features through the zero cross-attention block. However, StableVITON (Kim et al., 2023) aligns the global features and local features of clothing separately without considering direct interaction between the two, making it difficult to refine the details while preserving global semantic information. Therefore, we introduce the joint attention blocks, building upon the foundation of maintaining two encoders to extract the global and local features of clothing, respectively. We first align the global and local features to the target person, then selectively fuse the two to enhance the preservation of local texture and global semantic information in the original clothing.
3. method
While most existing virtual try-on methods are designed solely for frontal-view scenarios, we present a novel approach to handle both frontal-view and multi-view virtual try-on tasks, along with a multi-view virtual try-on dataset MVG comprising try-on images captured from five different views. Examples of our dataset are shown in Figure 2(b). Formally, given a person image $x\in R^{H\times W\times 3}$ in a frontal, side, or back view, along with a frontal view clothing $c_{f}\in R^{H\times W\times 3}$ and a back view clothing $c_{b}\in R^{H\times W\times 3}$, our goal is to generate the result of the person wearing the clothing in its view. Considering the substantial differences between the front and back of most clothing items, compared to previous frontal-view virtual try-on methods, another challenge we face is how to make informed decisions regarding the two provided clothing images based on the target person’s pose, ensuring a natural try-on result across multiple views.
In this work, we use an image inpainting diffusion model (Yang et al., 2023) as the baseline of our method, it concatenates the noisy image $z_{t}$, the encoded clothing-agnostic image $\mathcal{E}(a)$, and the resized clothing-agnostic mask $m$ in the channel dimension and feeds them into the backbone as spatial input. Therein, $a$ is the masked person image $x$ using the inpainting mask $M\in R^{H\times W\times 1}$. In addition, we use an existing method to initially deform the clothing and paste it on $a$. $m$ is the result of resizing $M$ to latent space and $\mathcal{E}$ represents VAE encoder (Kingma and Welling, 2013) in latent diffusion model (Rombach et al., 2022). While utilizing CLIP image encoder (Radford et al., 2021) to encode clothing as the global condition of the diffusion model, we also introduce an additional encoder (Yang et al., 2023; Zhang et al., 2023) to encode clothing to provide more refined local conditions. Since both the frontal and back view clothing need to be encoded, directly sending both into the backbone as conditions may result in confusion of clothing features. To alleviate this problem, we propose a view-adaptive selection mechanism. Based on the similarity between poses of the person and two clothes, it conducts hard-selection when extracting global features and soft-selection when extracting local features. To enhance high-frequency details in global features using local ones, while preserving semantic information in clothing, we introduce a joint attention block. It first independently aligns global and local features to the person ones and then selectively fuses them. Figure 3(a) depicts an overview of our proposed method.
Figure 2. Comparison between previous datasets and our proposed MVG dataset. (a) is the datasets used by the previous work, which only have clothing and person in the frontal-view. In contrast, each sample in our dataset (b) offers images from five different views. By utilizing the first and fifth images to provide frontal and back view clothing, we can reconstruct multi-view virtual try-on results.
[…]
[Figure 2. Fully described in the text.]
Figure 3. (a) Overview of MV-VTON. It encodes frontal and back view clothing into global features using the CLIP image encoder (Radford et al., 2021) and extracts multi-scale local features through backbone encoder’s trainable copy $\mathcal{E}_{l}$. Both features act as conditional inputs for the decoder of backbone. Besides, both features are selectively extracted through view-adaptive selection mechanism. (b) Soft-selection modulates the clothing features on frontal and back view, respectively, based on the similarity between the clothing’s pose and the person’s pose. Then the features from both views are concatenated in the channel dimension.
[…]
[Figure 3. Fully described in the text.]
3.1. Preliminaries for Diffusion Models
Diffusion Models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022) have demonstrated strong capabilities in visual generation, which transforms a Gaussian distribution into a target distribution by iterative denoising. In particular, Stable Diffusion (Rombach et al., 2022) is a widely used generative diffusion model, which consists of a CLIP text encoder  (Radford et al., 2021) $\mathcal{E}_{T}$, a VAE (Kingma and Welling, 2013) encoder $\mathcal{E}$ as well as decoder $\mathcal{D}$, and a time-conditional denoising model $\epsilon_{\theta}$. The text encoder $\mathcal{E}_{T}$ encodes the input text prompt $y$ as conditional input. The VAE encoder $\mathcal{E}$ compresses the input image $I$ into latent space to get the latent variable $z_{0}=\mathcal{E}(I)$. In contrast, the VAE decoder $\mathcal{D}$ decodes the output of backbone from latent space to pixel space. Through the VAE encoder $\mathcal{E}$, at an arbitrary time step t, the forward process is performed:
(1) $\alpha:={\textstyle\prod_{s=1}^{t}(1-\beta_{s})}~{},\indent z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$
where $\epsilon\sim\mathcal{N}(0,1)$ is the random Gaussian noise and $\beta$ is a predefined variance schedule. The training objective of the diffusion model is to acquire a noise prediction network that minimizes the disparity between the predicted noise and the noise added to ground truth. The loss function can be defined as,
(2) $\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),y,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{T}(y))\right\|_{2}^{2}],$
where $t$ is the time step, and $z_{t}$ represents the encoded image $\mathcal{E}(I)$ with random Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$ added.
In our work, we use an exemplar-based inpainting model (Yang et al., 2023) as a backbone, which employs an image $c$ rather than texts as the prompt and then encode $c$ by the image encoder $\mathcal{E}_{I}$ of CLIP (Radford et al., 2021). Thus, the loss function in Eq. (2) can be modified as,
(3) $\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),c,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{I}(c))\right\|_{2}^{2}].$
3.2. View-Adaptive Selection
For multi-view virtual try-on task, given the substantial differences between the frontal and back view, as illustrated in Figure 2(b), it’s imperative to extract and assign the features of frontal and back view clothing for the person tendentiously. Actually, based on the pose of the target person, we can determine which view’s clothing should be given more attention during the try-on process. For example, if the target pose resembles the pose in the fourth column of Figure 2(b), it’s evident that we should rely more on the characteristics of the back view clothing to generate the try-on result. Specifically, we propose a view-adaptive selection mechanism to achieve this purpose, including hard- and soft-selection.
Hard-Selection for Global Clothing Features. The image encoder of CLIP (Radford et al., 2021) is used to extract global features of clothing. Before this process, we perform hard-selection on the clothing images under the frontal and back view based on the similarity between the garments’ pose and the person’s pose. It means that we only select one piece of clothing that is closest to the person’s pose as the input of the image encoder, since it is enough to cover global semantic information. When generating pre-warped clothing in $\mathcal{E}(a)$, the selection is also performed. Implementation details of hard-selection can be found in the supplementary material.
Soft-Selection for Local Clothing Features. The multi-scale local features of frontal and back view clothing are extracted through an additional encoder $\mathcal{E}_{l}$. Considering that relying solely on the clothing from the frontal or back view may not be sufficient when reconstructing the try-on results under certain specific scenes, such as the third column shown in Figure 2(b). In these cases, it may be necessary to incorporate clothing features from both the frontal and back views. However, simply combining the two may lead to confusion of features. Instead, we introduce soft-selection block to modulate their local features, respectively, as shown in Figure 3(b). Firstly, the person’s pose $p_{h}$, frontal-view clothing’s pose $p_{f}$ and back view clothing’s pose $p_{b}$ are encoded by the pose encoder $\mathcal{E}_{p}$ to obtain their respective features $\mathcal{E}_{p}(p_{h})$, $\mathcal{E}_{p}(p_{f})$ and $\mathcal{E}_{p}(p_{b})$. Details of the pose encoder can be found in the supplementary material. When processing frontal-view clothing, in $i$-th soft-selection block, we map $\mathcal{E}_{p}(p_{h})$ and $\mathcal{E}_{p}(p_{f})$ to $P^{i}_{h}$ and $P^{i}_{f}$ through a linear layer with weights $W_{h}^{i}$ and $W_{f}^{i}$, respectively. For the features of frontal-view clothing $c_{f}^{i}$, it will be mapped to $C^{i}_{f}$ through a linear layer with weights $W_{c}^{i}$. Then we calculate the similarity between the person’s pose and frontal-view clothing’s pose to get the selection weights of frontal-view clothing, i.e.,
(4) $weights=softmax(\frac{P_{h}^{i}(P_{f}^{i})^{T}}{\sqrt{d}}),$
where $weights$ represents the selection weights of frontal-view clothing, and $d$ represents the dimension of these matrices. Assuming that the person’s pose is biased towards the front, as depicted in the second column of Figure 2(b), the similarity between the person’s pose and the front view clothing’s pose will be higher. Consequently, the corresponding clothing features will be enhanced by $weights$, and vice versa.
The features of back view clothing $c_{b}^{i}$ undergo similar processing. Consequently, one of the features $c_{f}^{i}$ and $c_{b}^{i}$ in the two views will be enhanced while the other will be weakened. Finally, the two selected clothing features are concatenated along the channel dimension as the local condition $c_{l}^{i}$ of backbone.
Figure 4. Overview of the proposed joint attention block.
[…]
[Figure 4. Fully described in the text.]
3.3. Joint Attention Block
Global clothing features $c_{g}$ provide identical conditions for blocks at each scale of U-Net, and multi-scale local clothing features $c_{l}$ allow for reconstructing more accurate details. We present a joint attention block to align $c_{g}$ and $c_{l}$ with the current person features, as shown in Figure 4. Additionally, to retain most of the semantic information in global features $c_{g}$, we use local features $c_{l}$ to refine some lost and erroneous detailed texture information in $c_{g}$ by selective fusion.
Specifically, in the $i$-th joint attention block, we first calculate self-attention for the current features $f_{in}^{i}$, followed by double cross-attention where the queries (Q) come from $f_{in}^{i}$ and global features $c_{g}$ serve as one set of keys (K) and values (V), while local features $c_{l}^{i}$ serve as another set of keys (K) and values (V).
After aligning to the person’s pose through cross-attention, the clothing features $c_{g}$ and $c_{l}^{i}$ are selectively fused in channel-wise dimension, i.e.,
(5) $f_{out}^{i}=softmax(\frac{Q^{i}_{g}(K^{i}_{g})^{T}}{\sqrt{d}})V^{i}_{g}+\lambda\odot softmax(\frac{Q^{i}_{l}(K^{i}_{l})^{T}}{\sqrt{d}})V^{i}_{l}~{},$
where $Q^{i}_{g},K^{i}_{g},V^{i}_{g}$ represent the Q, K, V of global branch, $Q^{i}_{l},K^{i}_{l},V^{i}_{l}$ represent the Q, K, V of local branch, $\lambda$ is the learnable fusion vector, $\odot$ represents channel-wise multiplication, and $f_{out}^{i}$ represents the clothing features after selective fusion. By engaging and fusing the global and local features of garments, we can enhance the retention of high-frequency details like texts and patterns in original clothing.
3.4. Training Objectives
As stated in Section 3.1, diffusion models learn to generate images from random Gaussian noise. However, the training objective in Eq. (3) is performed in latent space, and does not explicitly constrain the generated results in visible image space, resulting in slight differences in color from the ground truth. To alleviate the problem, we additionally employ $\ell_{1}$ loss $\mathcal{L}_{1}$ and perceptual loss (Johnson et al., 2016) $\mathcal{L}_{perc}$. The $\mathcal{L}_{1}$ loss is calculated by
(6) $\mathcal{L}_{1}=\left\|\hat{x}-x\right\|_{1},$
where $\hat{x}$ is the reconstructed image using Eq. (1). The perceptual loss is calculated as,
(7) $\mathcal{L}_{perc}=\sum_{k=1}^{5}\left\|\phi_{k}(\hat{x})-\phi_{k}(x)\right\|_{1},$
where $\phi_{k}$ represents the $k$-th layer of VGG (Simonyan and Zisserman, 2014). Totally, the overall training objective can be written as,
(8) $\mathcal{L}=\mathcal{L}_{LDM}+\lambda_{1}\mathcal{L}_{1}+\lambda_{perc}\mathcal{L}_{perc}~{},$
where $\lambda_{1}$ and $\lambda_{perc}$ are the balancing weights.
Table 1. Quantitative comparison with previous work on paired setting. For multi-view virtual try-on task, we show results on our proposed MVG dataset. For frontal-view virtual try-on task, we show results on VITON-HD dataset (Choi et al., 2021) and DressCode dataset (Morelli et al., 2022). The best results have been bolded. Note that all previous works have been finetuned on our proposed MVG dataset when comparing on multi-view virtual try-on task.
Method MVG VITON-HD DressCode - Upper Body
LPIPS↓ SSIM↑ FID↓ KID↓ LPIPS↓ SSIM↑ FID↓ KID↓ LPIPS↓ SSIM↑ FID↓ KID↓
Paint by Example (Yang et al., 2023) 0.120 0.880 54.38 14.95 0.150 0.843 13.78 4.48 0.078 0.899 15.21 4.51
PF-AFN (Ge et al., 2021b) 0.139 0.873 49.47 12.81 0.141 0.855 7.76 4.19 0.091 0.902 13.11 6.29
GP-VTON (Xie et al., 2023) - - - - 0.085 0.889 6.25 0.77 0.236 0.781 19.37 8.07
LaDI-VTON (Morelli et al., 2023) 0.069 0.921 29.14 4.39 0.094 0.872 7.08 1.49 0.063 0.922 11.85 3.20
DCI-VTON (Gou et al., 2023) 0.062 0.929 25.71 0.95 0.074 0.893 5.52 0.57 0.043 0.937 11.87 1.91
StableVITON (Kim et al., 2023) 0.063 0.929 23.52 0.46 0.073 0.888 6.15 1.34 0.040 0.937 10.18 1.70
Ours 0.050 0.936 22.17 0.36 0.068 0.897 5.42 0.48 0.040 0.941 8.26 1.39
Table 2. Unpaired setting’s quantitative results on our MVG dataset and VITON-HD dataset (Choi et al., 2021). The best results have been bolded.
Method MVG VITON-HD
${\rm FID_{u}}$ ${\rm KID_{u}}$ ${\rm FID_{u}}$ ${\rm KID_{u}}$
Paint by Example (Yang et al., 2023) 43.79 5.92 17.27 4.56
PF-AFN (Ge et al., 2021b) 47.38 7.04 21.18 6.57
GP-VTON (Xie et al., 2023) - - 9.11 1.21
LaDI-VTON (Morelli et al., 2023) 36.61 3.39 9.55 1.83
DCI-VTON (Gou et al., 2023) 36.03 3.79 8.93 1.07
StableVITON (Kim et al., 2023) 35.85 4.22 9.86 1.09
Ours 34.48 2.92 8.67 0.78
4. experiments
4.1. Experiments Setting
Multi-View Dataset: For the proposed multi-view virtual try-on task, we collect MVG dataset containing 1,009 samples. Each sample contains five images of the same person wearing the same garment from five different views, for a total of 5,045 images, as shown in Figure 2(b). The image resolution is about 1K. Collection details of this dataset are provided in the supplementary material. In this work, we split it into a training set with 812 samples and a test set with 197 samples. During training, we extract the clothing from the frontal and back view images (the first and fifth columns of Figure 2(b)) to reconstruct the try-on results from other views (the second to fourth columns of Figure 2(b)), and use the real images from the corresponding views as ground truth. During testing, users only need to provide two images in frontal and back views to generate try-on results of a person from multiple views.
Frontal-View Datasets. The proposed method can also be applied to frontal-view virtual try-on task. Our frontal-view experiments are carried out on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. VITON-HD (Choi et al., 2021) has 13,679 frontal-view person and upper-body clothing image pairs with a resolution of 1,024×768. DressCode (Morelli et al., 2022) contains three categories: upper-body clothing, lower-body clothing, and dresses, with a total of more than 53,000 image pairs of frontal-view clothing and persons. Among them, upper-body clothing contains 15,363 samples. For the division of the datasets, we follow the previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023). For VITON-HD (Choi et al., 2021) dataset, we use 2,032 pairs as the test set and 11,647 pairs as the training set. For DressCode (Morelli et al., 2022) dataset, we employ 1,800 pairs as the test set, and the rest as the training set.
Evaluation Metrics. Following previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023; Ge et al., 2021b; Xie et al., 2023; Yang et al., 2023), we use four metrics to evaluate the performance of our method: Structural Similarity (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Frechet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018). LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004) tend to measure the consistency of the generated try-on results with ground truth. FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) are employed to measure the realism of images. Specifically, for paired test setting, which means directly using the paired data in the dataset, we utilize the above four metrics for evaluation. For unpaired test setting, which means that the given garment is different from the garment originally worn by target person, we use FID and KID for evaluation, and in order to distinguish them from the paired setting, we named them ${\rm FID_{u}}$ and ${\rm KID_{u}}$ respectively.
Figure 5. Qualitative comparison on multi-view virtual try-on task.
[…]
[Figure 5. Fully described in the text.]
Figure 6. Qualitative comparison on frontal-view virtual try-on task.
[…]
[Figure 6. Fully described in the text.]
Implementation Details. We use Paint by Example (Yang et al., 2023) as the backbone of our method and copy the weights of its encoder to initialize $\mathcal{E}_{l}$. The hyper-parameter $\lambda_{1}$ is set to 1e-1, and $\lambda_{perc}$ is set to 1e-4. We train our model on 2 NVIDIA Tesla A100 GPUs for 40 epochs with a batch size of 4 and a learning rate of 1e-5. We use AdamW (Loshchilov and Hutter, 2017) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$.
Comparison Settings. We compare our method with Paint By Example (Yang et al., 2023), PF-AFN (Ge et al., 2021b), GP-VTON (Xie et al., 2023), LaDI-VTON (Morelli et al., 2023), DCI-VTON (Gou et al., 2023) and StableVITON (Kim et al., 2023) on both frontal-view and multi-view virtual try-on tasks. For multi-view virtual try-on, we compare these methods on the proposed MVG dataset. For the sake of fairness, we fine-tune the previous methods on the MVG dataset according to its original training settings. Since previous methods can only input a single clothing image, we input frontal and back view clothing respectively and select the best result. For frontal-view virtual try-on, we compare these methods on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. Following previous works’ settings, the proposed MV-VTON only inputs one frontal-view garment during training and inference. When evaluating previous methods, we directly use their released models to test, and train the models if not available.
4.2. Quantitative Evaluation
Table 1 reports the quantitative results on the paired setting, and Table 2 shows the unpaired setting’s results. On the multi-view virtual try-on task, as can be seen, thanks to the view-adaptive selection mechanism, our method can reasonably select clothing features according to the person’s pose, so it is better than existing methods in various metrics, especially on LPIPS and SSIM. Furthermore, owing to the joint attention block, our approach excels in preserving high-frequency details of the original garments across both frontal-view and multi-view virtual try-on scenarios, thus achieving superior performance in these metrics.
4.3. Qualitative Evaluation
Multi-View Virtual Try-On. As shown in Figure 5, MV-VTON generates more realistic multi-view results compared to the previous five methods. Specifically, in the first row of Figure 5, due to the lack of adaptive selection of clothes, previous methods have difficulty in generating hoods of the original cloth. Moreover, in the second and third rows, previous methods often struggle to maintain fidelity to the original garments. In contrast, our method effectively addresses the aforementioned problems and generates high-fidelity results.
Frontal-View Virtual Try-On. As shown in Figure 6, our method also demonstrates superior performance over existing methods on frontal-view virtual try-on task, particularly in retaining clothing details. Specifically, in the second row of the Figure 6, our method better preserves the literal ’KORS’ in the clothing. In addition, benefiting from the selective fusion of global and local clothing features in joint attention blocks, our method not only preserves the shape of special garments (in the first row), but also faithfully generates complex patterns (in the third row).
Figure 7. Visual comparison of effect of the proposed view-adaptive selection mechanism on MVG dataset.
[…]
[Figure 7. Fully described in the text.]
Figure 8. Visual comparison of effect of the proposed joint attention block on VITON-HD (Choi et al., 2021) dataset.
[…]
[Figure 8. Fully described in the text.]
4.4. User Study
We conduct a user study involving 28 participants, each of which is presented with test results generated by our and previous methods. The participants evaluate the results based on fidelity (indicating whether the generated clothing is faithful to original clothing) and identity (indicating whether human body details such as the face and hands are preserved well) and choose the best one among all methods. We then calculate the proportion of the best images to the total results for each method. As shown in Table 3, the proposed method outperforms previous methods both on fidelity and identity.
4.5. Ablation Study
Effect of View-Adaptive Selection. We investigate the effect of view-adaptive selection on the multi-view virtual try-on task. Specifically, on the basis of no clothing selection, we add hard-selection when extracting global features and soft-selection when extracting local features, respectively. No hard-selection represents that we directly concatenate two garments’ features encoded by CLIP (Radford et al., 2021), and no soft-selection means that two clothing features are concatenated without passing soft-selection blocks. Comparison results are shown in Table 4 and Figure 7. As can be seen, without any selection, the performance is greatly reduced. Merely employing hard-selection or soft-selection can enhance performance, but the generated results still exhibit some unnaturalness, such as the unrealistic hoods and black stripes, as shown in Figure 7. In contrast, by combining hard-selection and soft-selection, our method can generate more faithful try-on results.
Table 3. User study results. We make comparisons with previous works on MVG and VITON-HD (Choi et al., 2021) datasets.
Method MVG VITON-HD
Fidelity↑ Identity↑ Fidelity↑ Identity↑
Paint by Example (Yang et al., 2023) 0.76% 0.33% 0.32% 1.66%
PF-AFN (Ge et al., 2021b) 0.37% 0.63% 0.17% 0.19%
GP-VTON (Xie et al., 2023) - - 11.4% 0.42%
LaDI-VTON (Morelli et al., 2023) 11.38% 19.81% 3.37% 9.57%
DCI-VTON (Gou et al., 2023) 3.77% 8.3% 14.13% 13.63%
StableVITON (Kim et al., 2023) 15.13% 29.2% 17.9% 27.42%
Ours 68.59% 41.73% 52.71% 47.11%
Table 4. Ablation study of our proposed view-adaptive selection mechanism on MVG dataset.
Hard Soft LPIPS↓ SSIM↑ FID↓ KID↓ ${\rm FID_{u}}$ ${\rm KID_{u}}$
$\times$ $\times$ 0.068 0.925 25.13 0.77 35.28 3.24
$\times$ $\surd$ 0.064 0.928 24.58 0.62 34.67 3.05
$\surd$ $\times$ 0.052 0.934 22.18 0.43 33.47 2.74
$\surd$ $\surd$ 0.050 0.936 22.18 0.35 33.44 2.69
Table 5. Ablation study of the joint attention block on MVG and VITON-HD (Choi et al., 2021) datasets.
Global Local LPIPS↓ SSIM↑ FID↓ KID↓ ${\rm FID_{u}}$ ${\rm KID_{u}}$
MVG $\surd$ $\times$ 0.062 0.929 25.71 0.95 36.01 3.78
$\times$ $\surd$ 0.058 0.931 26.16 1.21 36.29 3.91
$\surd$ $\surd$ 0.050 0.936 22.18 0.35 33.44 2.69
VITON-HD $\surd$ $\times$ 0.074 0.893 5.52 0.57 8.93 1.07
$\times$ $\surd$ 0.070 0.896 5.76 0.81 9.15 1.09
$\surd$ $\surd$ 0.069 0.897 5.43 0.49 8.67 0.78
Effect of Joint Attention Block. In order to demonstrate the effectiveness of fusing global and local features through joint attention blocks, we discard the global feature extraction branch and the local feature extraction branch respectively. Results are shown in Table 5 and Figure 8. As can be seen, relying solely on global features may lead to loss of details, such as the distorted text ’VANS’ in the first row and the missing letter ’C’ in the second row. Moreover, if only local features are provided, the results may also have unfaithful textures, such as artifacts on the person’s chest. Compared to them, we fuse global and local features through joint attention blocks, which can refine details in garments while preserving semantic information.
5. conclusion
We introduce a novel and practical Multi-View Virtual Try-ON (MV-VTON) task, which aims at using the frontal and back clothing to reconstruct the dressing results of a person from multiple views. To achieve the task, we propose a diffusion-based method. Specifically, the view-adaptive selection mechanism exacts more reasonable clothing features based on the similarity between the poses of a person and two clothes. The joint attention block aligns the global and local features of the selected clothing to the target person, and fuse them. In addition, we collect a multi-view garment dataset for this task. Extensive experiments demonstrate that the proposed method achieves state-of-the-art performance both on frontal-view and multi-view virtual try-on tasks, compared with existing methods.
Appendix
Appendix A IMPLEMENTATION DETAILS
Hard-Selection. In this section, we present more details about the proposed hard-selection for global clothing features. Specifically, in multi-view virtual try-on task, we use OpenPose (Cao et al., 2017; Simon et al., 2017; Wei et al., 2016) to extract the skeleton images of target person, frontal clothing and back clothing as pose information $p_{h}$, $p_{f}$, and $p_{b}$, respectively. After that, we decide whether to use frontal-view clothing or back-view clothing based on the relative positions of the target person’s left arm and right arm in the skeleton images. As shown in Figure A, if the right arm appears positioned to the left of the left arm in the skeleton image (columns one to three in Figure A), frontal-view clothing is chosen; otherwise, back-view clothing is preferred (columns four to five in Figure A). In addition, following previous works (Gou et al., 2023; Xie et al., 2023; Morelli et al., 2023), we adopt PF-AFN (Ge et al., 2021b) and GP-VTON (Xie et al., 2023) to obtain the pre-warped cloth on frontal-view try-on task; and we utilize StableVITON (Kim et al., 2023) to get it on multi-view try-on task.
Pose Encoder. The pose encoder is used to extract features of skeleton images. It is a tiny network that contains three blocks, followed by the layer normalization (Ba et al., 2016). Each block comprises one convolution layer, one GELU (Hendrycks and Gimpel, 2016) activation layer, and a down-sampling operation. We utilize the acquired pose embeddings as input for the proposed soft-selection block.
MVG Dataset. To construct dataset for multi-view virtual try-on (MV-VTON), we first collect a large number of videos from YOOX NET-A-PORTER [^1] https://net-a-porter.com, Taobao [^2] https://taobao.com and TikTok [^3] https://douyin.com websites, then filter out 1009 videos where the person in the video turns at least 180 degrees while wearing clothes. Afterwards, we divide each video into frames and handpick 5 frames to constitute a sample within our MVG dataset. Across these 5 frames, the person is captured from various angles, approximately spanning 0 (i.e., frontal view), 45, 90, 135, and 180 (i.e., back view) degrees, as shown in the first row of Figure B. In addition, following the previous DressCode dataset (Morelli et al., 2022), we employ SCHP model (Li et al., 2020) to extract the corresponding human parsing maps and utilize DensePose (Güler et al., 2018) to obtain the person’s dense labels, as shown in the second and third row of Figure B. Human parsing maps can be utilized to generate cloth-agnostic person images, which are necessary for training and inference processes.
Appendix B MORE QUALITATIVE RESULTS
Results on Multi-View VTON. In this section, we present more qualitative results on the MVG dataset. Specifically, as shown in Figure D, we show multiple groups of try-on results for the same person under different views, using the proposed method. In Figure D, the first column displays frontal-view and back-view garments, the second to fourth columns depict persons from different views, while the fifth to seventh columns showcase the corresponding try-on results. As can be seen, our method can generate realistic dressing-up results of the multi-view person from the given two views of clothing. Furthermore, our method can retain the details well on the original clothing (e.g., the buttons in the fifth row) and generate high-fidelity try-on images even under occlusion (e.g., hair occlusion in the second row). In conclusion, the proposed method exhibits outstanding performance on multi-view virtual try-on task.
Figure A. Visualization of the person and corresponding poses. We select one of garments based on the relative positions of left and right arms in the skeleton image when performing hard-selection on the multi-view virtual try-on task.
[…]
[Figure 1. Fully described in the text.]
Figure B. Examples of human parsing maps and dense pose in our dataset. The parsing maps can be used to synthesize cloth-agnostic person images.
[…]
[Figure 2. Fully described in the text.]
Results on Frontal-View VTON. In this section, we show more visual comparison results on VITON-HD (Choi et al., 2021) dataset and DressCode (Morelli et al., 2022) dataset. Previous works include Paint By Example (Yang et al., 2023), PF-AFN (Ge et al., 2021b), GP-VTON (Xie et al., 2023), LaDI-VTON (Morelli et al., 2023), DCI-VTON (Gou et al., 2023) and StableVITON (Kim et al., 2023). The results are shown in Figure E. In the first and second row of Figure E, it can be seen that our method better preserves the shape of the original clothing (e.g., the cuff in the second row), compared to the previous methods. In addition, our method outperforms previous methods in preserving high-frequency details, such as patterns on clothing in the fourth and sixth rows. Moreover, in contrast to previous methods, MV-VTON is not constrained by specific types of clothing and can achieve highly realistic effects across a wide range of garment styles (e.g., the garment in the third row and the collar in the eighth row). In summary, our method also has superiority on frontal-view virtual try-on task.
High Resolution Results on Frontal-View VTON. Following the previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023), the above results are all obtained at 512$\times$384 resolution. In this section, we present more results at 1024$\times$768 resolution on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets, as shown in the Figure F. Specifically, we utilize the model trained at 512$\times$384 resolution to directly test at 1024$\times$768 resolution. Despite the difference in resolutions between training and testing, our method can also produce high-fidelity try-on results. For instance, the generated images can preserve both the intricate patterns and text adorning the clothing (in the first row) while also effectively maintaining their original shapes (in the last row).
Appendix C LIMITATIONS
Despite outperforming previous methods on both frontal-view and multi-view virtual try-on tasks, our method does not perform well in all cases. Figure C displays some unsatisfactory try-on results. As can be seen, although our method can preserve the shape and texture of original clothing (e.g., the ’DIESEL’ text in the first row), it is difficult for it to fully preserve some smaller or more complex details (e.g., the parts circled in red). The reason for this phenomenon may be that these details are easily lost when inpainting in latent space. We will try to solve this issue in future work.
Figure C. Visualization of bad cases on VITON-HD dataset.
[…]
[Figure 3. Fully described in the text.]
Figure D. Qualitative results on multi-view virtual try-on task.
[…]
[Figure 4. Fully described in the text.]
Figure E. Qualitative comparison on frontal-view virtual try-on task.
[…]
[Figure 5. Fully described in the text.]
Figure F. Qualitative results of 1024$\times$768 resolution on frontal-view virtual try-on task.
[…]
[Figure 6. Fully described in the text.]